summaryrefslogtreecommitdiff
path: root/MeetingNotes25Jan04.mdwn
blob: 633467580e88560dd965b03cd47c2fd39fdd9eff (plain)
1
2
3
4
5
6
7
8
9
10
11
12
13
14
15
16
17
18
19
20
21
22
23
24
25
26
27
28
29
30
31
32
33
34
35
36
37
38
39
40
41
42
43
44
45
46
47
48
49
50
51
52
53
54
55
56
57
58
59
60
61
62
63
## <a name="Hybrid Meeting Minutes 01/25/04"></a> Hybrid Meeting Minutes 01/25/04

----

**In attendance:** Rainer Anacker, Peter Bold, Tim Brandon, Eric Hettinger, Maggie Emery, Brian O’Neel, Duftan Sandoz, Jim Ward, Evan Waymire

<div>
  <center>[[MeetingNotes25Jan04/pic012504B.jpg]]</center>
</div>

**Topics Discussed:**

We discussed the logistics of the next static test fire to be conducted on Jan 31st in Estacada. We will attempt to fire three test grains. One from our last paraffin pour and two that were poured today.

The purpose of the test will be to:

- Test the new oxygen shut-off system.
- Test the new 100 ft serial cable
- Test the effect of not having a nitrogen purge cycle on the paraffin
- Test the effect on the paraffin of moving the purge nitrogen to the other side of the metering orifice.
- Test the effect on the paraffin of reducing the ignition ring size.
- Test the effect on the paraffin of reducing the time from ignition ring start to oxygen flow initiation.
- Test the effect of extending the paraffin/graphite interface to minimize the change of premature liner or combustion chamber burn-through.
- Test a better method of securing the igniter to the ignition ring.

Some of the work done today was to pour the new paraffin fuel grains. The grains were weighed and labeled. We also changed over the plumbing to accommodate the new 4-way solenoid valve. This will hopefully remedy our oxygen shut-off problem. We discussed and worked on the problem of interfacing the 110 VAC inverter into the system.

When we finished these tasks we did a flow and switching test of the modified plumbing under pressure (80 psi nitrogen, 800 psi oxygen). The first attempt failed due to the new 100 ft. serial cable. Once the cable was switched out with the original 25’ cable the system worked fine.

We also reviewed the progress made on the new test stand, which will use an adjustable pitch I-beam and trolley configuration. The new test stand will be fully sensored so that we can start collecting data.

As an aside we also did a few ‘non-scientific’ experiments with some left over cryogenic liquid nitrogen and a few threaded stainless steel fittings to see if the extreme cold would cause shrinkage and leakage. A future liquid oxygen system may have this problem. The cryogenic nitrogen was not under pressure but showed no signs of leaking fro the threaded Swagelok connectors. The only thing that we can say for sure is that liquid nitrogen is VERY cold.

The next meeting will be at the static test fire location Saturday January 31st at noon.

**Action Items:**

- **Peter Bold**

1. Weld Analysis

- **Brian O’Neel:**

1. Pick up new oxygen bottle
2. Trim fuel gains
3. Insert ignition rings

- **Jim Ward:**

1. Look into 100 ft serial cable problem
2. Charge 12V battery

- **Evan Waymire**

1. Start researching funding and grants

<div>
  <center>[[MeetingNotes25Jan04/pic012504A.jpg]]</center>
</div>

<div>
  <center>[[MeetingNotes25Jan04/pic012504C.jpg]]</center>
</div>